Find the Maclaurin series for the function f(x) = square root of

(x2+2-x)5

1
Expert's answer
2022-04-29T13:14:19-0400

"p_0(0)=f(0)=\sqrt{2^5}=\sqrt{32}"

"f'(x)=\frac{5(2x-1)(x^2-x+2)^{3/2}}{2}"

"p_1(0)=f(0)+f'(0)x=\sqrt{32}-\frac{5\sqrt{8}}{2}x"

"f''(x)=5(x^2-x+2)^{3/2}+\frac{15(2x-1)^2\sqrt{x^2-x+2}}{4}"

"p_2=\sqrt{32}-\frac{5\sqrt{8}}{2}x+ \frac{5\sqrt{8}}{2}x^2+\frac{15\sqrt{2}}{8}x^2"